OpenStudy (kpop4life123):

FAST I NEED HELP!!!!!!!!!!!!!!! I WILL FAN AND MEDAL!!!!!!!!!!!! Which set of line segments could create a right triangle? (4 points) 15, 30, 35 15, 36, 39 15, 20, 29 5, 15, 30

OpenStudy (anonymous):

a^2 + b^2 = c^2. So your answer is going to be the first two numbers squared added together is going to equal the third number squared.

OpenStudy (kpop4life123):

THANK YOU! :)

OpenStudy (anonymous):

So if we try it out on one of your options, let's say option b, we need to substitute so the expression is 15^2 + 36^2 = 39^2. So now we need to solve the squares to get 225+1296=1521. Now all we need to do is find the square root of 1521 which is 39. So we know that b is the answer because all of the numbers substitute and equate properly.

OpenStudy (anonymous):

Does that make sense?